### v1

# A note on Combinatorial Invariance of Kazhdan–Lusztig polynomials

Abstract.
We introduce the concepts of an amazing hypercube decomposition and a double shortcut for it, and use these new ideas to formulate a conjecture implying the Combinatorial Invariance Conjecture of the Kazhdan–Lusztig polynomials for the symmetric group. This conjecture has the advantage of being combinatorial in nature.
Key words and phrases:
Kazhdan-Lusztig polynomials, Combinatorial invariance, hypercube decompositions
2020 Mathematics Subject Classification:
05E10 - 05E16 (primary), 20F55 (secondary)
1. Introduction
The Combinatorial Invariance Conjecture for Kazhdan–Lusztig polynomials was formulated independently by Lusztig and Dyer in the ’80’s. Its version for the symmetric group $W$ is the following.
Conjecture 1.1.
Let $u,v\in W$. The Kazhdan–Lusztig polynomial $P_{u,v}(q)$ (or, equivalently, the Kazhdan–Lusztig $\widetilde{R}$-polynomial $\widetilde{R}_{u,v}(q)$) depends only on the isomorphism class of the Bruhat interval $[u,v]$ as a poset.
It has been the focus of active research for the past forty years and has recently received important new inputs by [3] and [6], where the concept of a hypercube decomposition was introduced and used to give a conjectural formula for the Kazhdan–Lusztig polynomials of the symmetric group that implies Conjecture 1.1.
Successively, a conjectural formula for $\widetilde{R}$-polynomial $\widetilde{R}_{u,v}(q)$ of the symmetric group was given in terms of shortcuts ([5]) of a hypercube decomposition. Also this formula would imply Conjecture 1.1. A further work studying hypercube decompositions is [1], where the authors give another conjecture implying Conjecture 1.1 and prove combinatorial invariance for elementary intervals, which marks an improvement of the result in [4] for lower Bruhat intervals.
In this short note, we introduce the concepts of an amazing hypercube decomposition and a double shortcut for it, and use these new ideas to formulate yet another conjecture implying the conjecture in [5] and thus ultimately Conjecture 1.1. This new conjecture has the advantage of being combinatorial in nature, with no mention of Kazhdan–Lusztig polynomials: a purely graph theoretic property of the Bruhat interval would imply the Combinatorial Invariance Conjecture for the symmetric group.
The conjecture presented in this paper has been verified (in a strong form) up to $S_{6}$ by computer calculations.
2. Preliminaries
In this section we recall the basic definitions and set the notation to be used. We refer the reader to [2] for undefined terminology concerning Coxeter groups.
We recall the definition of a hypercube decomposition from [3]. Let $(W,S)$ be a Coxeter group of type $A$. We denote by $B(W)$ the Bruhat graph of $W$.
Given a set $E$, let ${\mathcal{P}}(E)$ denote the directed Boolean algebra on $E$, i.e. the directed graph having the power set of $E$ as vertex set and where $I\rightarrow J$ if $I$ is obtained from $J$ by removing one element.
Let $p\in W$ and $E$ be a set of arrows of $B(W)$ all having $p$ as target. Then $E$ spans a hypercube if there exists a unique embedding of directed graphs $\theta:{\mathcal{P}}(E)\rightarrow W$ sending the directed edge $E\setminus{\{\alpha\}}\rightarrow E$ to $\alpha$, for all $\alpha\in E$. Furthermore, $E$ spans a hypercube cluster if every subset $E^{\prime}$ of $E$ consisting of edges with pairwise incomparable sources (with respect to Bruhat order) spans a hypercube.
Let $u,v\in W$ and $z\in[u,v]$. We say that $z$ is an upper hypercube decomposition of $[u,v]$ provided that:
  1. (1)
    $[z,v]$ is diamond complete (with respect to $[u,v]$), meaning that, if there exist $x\in[u,v]$ and $a_{1},a_{2},y\in[z,v]$, $a_{1}\neq a_{2}$, such that $x\rightarrow a_{1}\rightarrow y$, $x\rightarrow a_{2}\rightarrow y$, then $x\in[z,v]$;
  2. (2)
    for all $p\in[z,v]$, the set $E^{p}=\{x\rightarrow p:x\notin[z,v]\}$ spans a hypercube cluster.
Similarly, taking the dual versions of the above definitions, we obtain the concept of a lower hypercube decomposition.
Following [5], we give the next definition. We denote by $\operatorname{d}(x,y)$ the distance function of $B(W)$, i.e., the minimum of the lengths of paths from $x$ to $y$.
Definition 2.1.
Let $u,v,z\in W$ with $z\in[u,v]$. We let
$W_{[u,v]}^{z}:=\{p\in[z,v]:\operatorname{supp}(\Gamma)\cap[z,v]=\{p\}\textrm{ for all paths $\Gamma$ from $u$ to $p$ of length $\operatorname{d}(u,p)$}\},$
and
$\widetilde{R}_{u,v}^{z}(q):=\sum_{p\in W_{[u,v]}^{z}}q^{\operatorname{d}(u,p)}\widetilde{R}_{p,v}(q).$
We call the elements in $W_{[u,v]}^{z}$ the (upper) shortcuts of $[u,v]$ with respect to $z$. Furthermore, we say that an element $z$ in $[u,v]$ is an (upper) $R$-element for $[u,v]$ if $\widetilde{R}_{u,v}^{z}=\widetilde{R}_{u,v}$.
Remark 2.2.
The elements $\min([u,v]\cap W_{S\setminus\{s_{n-1}\}}v)$, $\min([u,v]\cap W_{S\setminus\{s_{1}\}}v)$, $\min([u,v]\cap vW_{S\setminus\{s_{n-1}\}})$, and $\min([u,v]\cap vW_{S\setminus\{s_{1}\}}v)$ are always upper hypercube decompositions for the interval $[u,v]$ (possibly not mutually distinct). We call these elements the standard upper hypercube decompositions for $[u,v]$. Note that these hypercube decompositions are called canonical in [5].
Remark 2.3.
We note that the concept of an upper shortcut can be equivalently defined in the following way. Let $u,v\in W$, and $z$ be an upper hypercube decomposition of $[u,v]$. An element $p$ in $[z,v]$ is an upper shortcut of $[u,v]$ with respect to $z$ if $\operatorname{d}(u,p)<\operatorname{d}(u,x)$ for all $x$ in $[z,p]$ satisfying $\operatorname{d}(x,p)=1$.
3. Amazing hypercube decompositions and double shortcuts
Definition 3.1.
Let $u,v\in W$. We say that an upper hypercube decomposition $z$ of $[u,v]$ is amazing provided that, for all $x\in[u,v]$, the intersection $[z,v]\cap[x,v]$ has a minimum $z\wedge x$, and this minimum $z\wedge x$ is a (necessarily amazing) hypercube decomposition of $[x,v]$. Furthermore, we say that an amazing hypercube decomposition $z$ of $[u,v]$ is an (upper) amazing $R$-element for $[u,v]$ if $z\wedge x$ is an (upper) $R$-element for $[x,v]$ for all $x\in[u,v]$.
Remark 3.2.
Let $u,v\in W$ and $z$ be a standard hypercube decomposition of $[u,v]$. Since $z\wedge x$ exists and is a standard hypercube decomposition of $[x,v]$ (of the same kind), we have that $z$ is an amazing hypercube decomposition, and, by [5, Corollary 3.10], an amazing $R$-element.
The following conjecture is a weak version of Conjecture of [5], which has been verified up to $W=S_{6}$.
Conjecture 3.3.
Let $W$ be a Coxeter group of type $A$, $u,v\in W$, and $z\in[u,v]$ be an amazing upper hypercube decomposition. Then $z$ is an $R$-element, i.e.
$\widetilde{R}_{u,v}=\sum_{p\in W_{[u,v]}^{z}}q^{\operatorname{d}(u,p)}\widetilde{R}_{p,v}.$
Conjecture 3.3 implies the Combinatorial Invariance Conjecture.
Definition 3.4.
Let $u,v\in W$, and $z$ and $z^{\prime}$ be two amazing upper hypercube decomposition of $[u,v]$. We say that an element $b$ in $[z,v]\cap[z^{\prime},v]$ is a $(z,z^{\prime})$-double shortcut of $[u,v]$ if there exists $p\in W^{z}_{[u,v]}$ such that $b\in W^{z\wedge z^{\prime}}_{[p,v]}$. We denote by $DS(z,z^{\prime})$ the multiset
$\{(\operatorname{d}(u,p)+\operatorname{d}(p,b),b):p\in W^{z}_{[u,v]}\text{ and}b\in W^{z\wedge z^{\prime}}_{[p,v]}\}.$
Theorem 3.5.
Let $W$ be a Coxeter group of type $A$, $u,v\in W$, and $z,z^{\prime}\in[u,v]$ be two amazing upper hypercube decomposition. Suppose that
  1. (1)
    $z$ is an amazing $R$-element,
  2. (2)
    $z^{\prime}\wedge x$ is an (amazing) $R$-element for all $x\in[u,v]\setminus\{u\}$,
  3. (3)
    $DS(z,z^{\prime})=DS(z^{\prime},z)$ (as multisets).
Then, $z^{\prime}$ is an (amazing) $R$-element.
Proof.
We have:
$\widetilde{R}_{u,v}$ $=$ $\sum_{p\in W_{[u,v]}^{z}}q^{\operatorname{d}(u,p)}\widetilde{R}_{p,v}$
$=$ $\sum_{p\in W_{[u,v]}^{z}}q^{\operatorname{d}(u,p)}\big{(}\sum_{b\in W_{[p,v]}^{z^{\prime}\wedge p}}q^{\operatorname{d}(p,b)}\widetilde{R}_{b,v}\big{)}$
$=$ $\sum_{(a,b)\in DS(z,z^{\prime})}q^{a}\widetilde{R}_{b,v}$
$=$ $\sum_{(a,b)\in DS(z^{\prime},z)}q^{a}\widetilde{R}_{b,v}$
$=$ $\sum_{p^{\prime}\in W_{[u,v]}^{z^{\prime}}}q^{\operatorname{d}(u,p^{\prime})}\big{(}\sum_{b\in W_{[p^{\prime},v]}^{z\wedge p^{\prime}}}q^{\operatorname{d}(p^{\prime},b)}\widetilde{R}_{b,v}\big{)}$
$=$ $\sum_{p^{\prime}\in W_{[u,v]}^{z^{\prime}}}q^{\operatorname{d}(u,p^{\prime})}\widetilde{R}_{p^{\prime},v},$
where the first equation follows by (1), the second equation follows by (2), the third equation by the definition of the multiset $DS(z,z^{\prime})$, the fourth by (3), the fifth by the definition of the multiset $DS(z^{\prime},z)$, and the sixth equation follows by (1). ∎
Fix $u,v\in W$. Let us define an equivalence relation on the set of amazing hypercube decompositions of $[u,v]$ as the transitive closure of the relation $z\sim z^{\prime}$ if $DS(z,z^{\prime})=DS(z^{\prime},z)$.
Conjecture 3.6.
Let $W$ be a Coxeter group of type $A$. The above relation is trivial (i.e., it has only one equivalence class).
The following conjecture is a weakening of Conjecture 3.6.
Conjecture 3.7.
Let $W$ be a Coxeter group of type $A$. Every equivalence class of the above relation contains an amazing $R$-element.
Remark 3.8.
Up to $S_{6}$, we checked by computer that, in fact, a much stronger statement than Conjecture 3.7 holds: $DS(z,z^{\prime})=DS(z^{\prime},z)$ for all amazing hypercube decompositions of any interval.
Theorem 3.9.
Conjecture 3.7 implies Conjecture 3.3.
Proof.
Towards a contradiction, let $[u,v]$ be an interval of minimal length having an amazing hypercube decomposition $z$ thath fails to be an $R$-element. By hypothesis, there exists a sequence $(z_{0},\ldots,z_{r})$ of amazing hypercube decompositions such that $z_{0}=z$, and $z_{r}$ is an amazing $R$-element, and $DS(z_{i-1},z_{i})=DS(z_{i},z_{i-1})$ for each $i\in[r]$. One gets a contradiction by iteratively applying Theorem 3.5. ∎
Acknowledgments: We would like to thank Grant T. Barkley, Francesco Brenti, and Christian Gaetz for stimulating discussions we had during the Workshop “Bruhat order: recent developments and open problems”held at the University of Bologna from 15 April 2024 to 19 April 2024. The authors are members of the I.N.D.A.M. group GNSAGA.